### v2

# Coupling Theory, Optimal Transport, and Strassen’s Theorem Beyond Regular Orders

Abstract.
Many results in probability (most famously, Strassen’s theorem on stochastic domination), characterize some relationship between probability distributions in terms of the existence of a particular structured coupling between them. Optimal transport, and in particular Kantorovich duality, provides a framework for formally unifying these results, but the standard duality theory requires topological conditions that are not satisfied in some settings. In this work, we investigate the extent to which Kantorovich duality still provides meaningful connections between distributional relations and their coupling counterparts, in the topologically irregular setting. Towards this end, we show that Strassen’s theorem “nearly holds” for topologically irregular orders but that the full theorem admits counterexamples. The core of the proof is a novel technical result in optimal transport, which shows that the Kantorovich dual problem is well-behaved for optimal transport problems whose cost functions can be written as a non-increasing limit of lower semi-continuous functions.
Key words and phrases:
coupling; duality; equivalence relations; eventual domination; Kantorovich duality; optimal transport; partial orders; stochastic domination; Strassen’s theorem
2020 Mathematics Subject Classification:
60A10, 28A35, 90C46
This material is based upon work for which DR was supported by a research contract with Sandia National Laboratories, a U.S. Department of Energy multimission laboratory.
1. Introduction
1.1. Background
A coupling of Borel probability measures $\mu,\nu$ on a Polish space $S$ is a Borel probability measure on $S\times S$ satisfying $\pi(S\times\,\cdot\,)=\mu$ and $\pi(\,\cdot\,\times S)=\nu$. Equivalently, a coupling is the joint law of a pair of $S$-valued random variables $(X,Y)$ defined on a common probability space in such a way that distributions of $X$,$Y$ are $\mu$$\nu$, respectively. Coupling is one of the core techniques in probability theory, where it is used ubiquitously in essentially all sub-disciplines; see [7, 33, 23] for introductions. At the same time, the study of couplings is interesting in its own right and a great deal literature has addressed fundamental questions in coupling theory, particularly in the setting of stochastic processes [13, 24, 12, 31, 9, 15, 28, 6, 5, 3, 20, 16] but also more generally [11, 29, 17].
The richness of the theory and applications of couplings led Thorisson to state the “working hypothesis” in [32] that “each meaningful distributional relation should have a coupling counterpart”. Well-known examples include Strassen’s theorem on stochastic domination [22], characterizations of coalescence of stochastic processes [13, 24, 12], characterization of Palm measures of stationary point processes [21], Strassen’s theorem on convex order [30], and more. The working hypothesis can also be seen as a practical matter, as it allows one to develop methods for testing partially-identified statistical models wherein the joint distribution of two random variables is unobserved while their marginal distributions are observed [10, 8]. The goal of this paper, plainly, is to better understand the extent to which Thorisson’s working hypothesis is true.
As a concrete example, consider Strassen’s theorem on stochastic domination [30, 22]. It states that if $\preceq$ is a partial order on $S$ such that the set $R_{\preceq}:=\{(x,y)\in S\times S:x\preceq y\}$ is sufficiently regular then the following are equivalent:
  • (i)
    We have $\mu(A)\leq\nu(A)$ for all monotone Borel sets $A\subseteq S$.
  • (ii)
    There exists a coupling $\pi$ of $\mu$ and $\nu$ satisfying $\pi(R_{\preceq})=1$.
Here, a Borel set $A\subseteq S$ is called monotone if $x\in A$ and $x\preceq y$ imply $y\in A$. In terms of the working hypothesis, statement (i) is a distributional relation and statement (ii) is its coupling counterpart.
Another example, called simply strong duality in existing literature, comes from the recent works [17, 26], which state that if $\cong$ is an equivalence relation on $S$ such that the set $R_{\cong}:=\{(x,y)\in S\times S:x\cong y\}$ is regular enough, then the following are equivalent:
  • (i)
    We have $\mu(A)=\nu(A)$ for all invariant Borel sets $A\subseteq S$.
  • (ii)
    There exists a coupling $\pi$ of $\mu$ and $\nu$ satisfying $\pi(R_{\cong})=1$.
Here, a Borel set $A\subseteq S$ is called invariant if $x\cong y$ implies that we have $x\in A$ if and only if $y\in A$. As before, we cast this in terms of the working hypothesis by noting that statement (i) is a distributional relation and statement (ii) is its coupling counterpart.
1.2. Problem Statement
Presently, we propose that Thorisson’s working hypothesis should be understood as a statement about Kantorovich duality for optimal transport problems whose cost functions take only values zero and one. For a formal explanation of this, suppose that $\mu$ and $\nu$ are Borel probability measures on $S$, and that $R\subseteq S\times S$ is a general measurable relation. The question of whether there exists a coupling $\pi$ of $\mu$ and $\nu$ satisfying $\pi(R)=1$ leads one to naturally consider the optimal transport problem
(1.1) $\begin{cases}\textnormal{minimize}&\int_{S\times S}(1-\mathbf{1}_{R})\,\textnormal{d}\pi\\ \textnormal{over}&\textnormal{couplings }\pi\textnormal{ of }\mu,\nu.\end{cases}$
At the same time, general results on Kantorovich duality show that (1.1) is equivalent (in a suitable sense) to the problem
(1.2) $\begin{cases}\textnormal{maximize}&\int_{S}\phi\,\textnormal{d}\mu+\int_{S}\psi\,\textnormal{d}\nu\\ \textnormal{over}&\textnormal{bounded measurable }\phi,\psi:S\to\mathbb{R}\\ \textnormal{with}&\phi(x)+\psi(y)\leq 1-\mathbf{1}_{R}(x,y)\textnormal{ for all }x,y\in S.\end{cases}$
Thus, by considering the distributional statement (i) that problem (1.2) has optimal value zero, and the coupling counterpart (ii) that problem (1.1) has optimal value zero, we have realized Thorisson’s working hypothesis as a statement about the duality between (1.2) and (1.1).
Towards making a rigorous proof out of this formal argument, we consider the setting above and we apply the result [25, Corollary 2.3.9] to get
(1.3) $\sup_{\phi,\psi}\left(\int_{S}\phi\,\textnormal{d}\mu+\int_{S}\psi\,\textnormal{d}\nu\right)=\inf_{\pi}\int(1-\mathbf{1}_{R})\,\textnormal{d}\pi,$
where the domains of the supremum and infimum are specified above. In other words, problems (1.1) and (1.2) always have the same value. Completing the argument, then, reduces to answering the following questions:
  • When is problem (1.2) “meaningful”?
  • When does problem (1.1) admit minimizers?
One way of interpreting the term “meaningful” is that we require some condition on the (marginal) probabilities assigned by $\mu,\nu$ to some events $A$ in a specified collection $\mathcal{A}$ of Borel sets. In other words, we would like to compare $\mu,\nu$ by the values they assign to the same event $A\in\mathcal{A}$, meaning the Kantorovich problem (1.2) should have one dual variable rather than two, and that the dual variables should be indicator functions rather than general measurable functions.
In previous works, affirmative answers to the above questions generally rely on strict topological assumptions about $R$. For instance, for a partial order $\preceq$, a standard assumption for Strassen’s theorem to hold (see [30]) is that $R_{\preceq}$ is a closed subset of $S\times S$. Analogously, for an equivalence relation $\cong$, a sufficient condition for strong duality to hold (see [17]) is that $R_{\cong}$ is a closed subset of $S\times S$. In fact, the result of [22] shows that if $R\subseteq S\times S$ is any closed set corresponding to a reflexive transitive relation (sometimes called pre-order), then a form of Strassen’s theorem holds. In this sense, we may regard Thorisson’s working hypothesis as completely settled for closed sets corresponding to reflexive transitive relations.
Some important relations (from the point of view of probability theory) do not correspond to closed sets and yet similar coupling results are known. For example, the equivalence relation of eventual equality or tail equivalence of sequences or paths has been studied in [13, 24, 12] and is closely related to some aspects of probabilistic potential theory and Markov chain mixing times. Another example is the equivalence relation of germ equivalence which has been studied in [9, 17]. A motivating example for the present work, which has not been previously studied, is the relation of eventual domination or tail domination, say on $S=\mathbb{R}^{\mathbb{N}}$ or more generally on the space of sequences on any partially ordered set; concretely, let $\leq^{\ast}$ be the relation on $S$ whereby $\{x_{n}\}_{n\in\mathbb{N}}\leq^{\ast}\{y_{n}\}_{n\in\mathbb{N}}$ if and only if $x_{n}\leq y_{n}$ for all sufficiently large $n\in\mathbb{N}$. In terms of probability theory, a coupling result for eventual domination à la Thorisson would allow one to characterize the possibility of long-term comparisons of, say, a Markov chain started from different initializations or of different Markov processes started from the same initialization.
Let us briefly digress to make some basic observations about the connection between the working hypothesis and optimal transport, namely that many of the usual objects of interest in abstract optimal transport (see[34]) take on concrete and interpretable forms when the cost function is $c=1-\mathbf{1}_{R}$ for a suitable relation $R$ on $S$. For example, suppose that $\preceq$ is a partial order and that $R=R_{\preceq}$. Then, a function $\phi:S\to\mathbb{R}$ is $c$-convex if and only if it is (non-negative and) monotone, and the $c$-transform $\phi^{c}$ of $\phi$ is (the negative of) its smallest monotone majorant. Also, a Borel set $A\subseteq S$ is monotone if and only if $\mathbf{1}_{A}:S\to\mathbb{R}$ is a $c$-convex function, and the monotone closure $A^{\prime}$ of $A$ is characterized via $\mathbf{1}_{A}^{c}=-\mathbf{1}_{A^{\prime}}$. For another example, suppose that $\cong$ is an equivalence relation and that $R=R_{\cong}$. Then, a function $\phi$ is $c$-convex if and only if it is (non-negative and) invariant, and the $c$-transform $\phi^{c}$ is (the negative of) its smallest invariant majorant. Also, a Borel set $A\subseteq S$ is invariant if and only if $\mathbf{1}_{A}$ is a $c$-convex function, and the saturation $A^{\prime}$ of $A$ is characterized via $\mathbf{1}_{A}^{c}=-\mathbf{1}_{A^{\prime}}$. Some of these observations have been made implicitly in the work [27], which provides a proof of Strassen’s theorem via optimal transport for partial orders $\preceq$ such that $R_{\preceq}$ is closed in $S\times S$.
Returning to the present work, our goal is investigate Thorisson’s working hypothesis for reflexive, transitive relations that need not be closed sets, but can be written as a countable increasing union of closed sets; that is, we focus on reflexive, transitive, $F_{\sigma}$ relations. As we have discussed, we have three motivations for this level of generality. First is the prior literature [17, 26] studying an equivalence relation $\cong$, which show that strong duality holds under some purely measure-theoretic conditions on $\cong$; notably, these works show that strong duality holds in some settings where standard approaches based on Kantorovich duality [27] can fail. Second is the natural question of whether Strassen’s theorem for closed partial order extends to $F_{\sigma}$ partial orders. Third is providing new results for the problem of eventual domination, which is reflexive and transitive but not antisymmetric (unless passing to a suitable quotient).
1.3. Main Results
Now we can state the results of the paper. Throughout, we use the notation that $\mathcal{B}(S)$ denotes the collection of Borel sets of $S$, and $\Pi(\mu,\nu)$ denotes the space of all couplings of Borel probability measures $\mu,\nu$. Our main result is the following, which states that (1.2) is “meaningful” but that (1.1) need not admit solutions:
Theorem 1.1.
If $S$ is a Polish space with its Borel $\sigma$-algebra $\mathcal{B}(S)$, and $R\subseteq S\times S$ is a reflexive, transitive, $F_{\sigma}$ relation, then we have
$\max_{A\in R^{\ast}}(\mu(A)-\nu(A))=\inf_{\pi\in\Pi(\mu,\nu)}(1-\pi(R))$
for any Borel probability measures $\mu,\nu$ on $S$, where
$R^{\ast}:=\{A\in\mathcal{B}(S):\textnormal{ if }x\in A\textnormal{ and }(x,y)\in R,\textnormal{ then }y\in A\}.$
The dual problem always admits optimizers, but the primal problem may not.
This result states that Thorisson’s working hypothesis “nearly holds” for reflexive, transitive, $F_{\sigma}$ relations. More precisely, it implies that the following distributional relation and its coupling counterpart are equivalent:
  • (i)
    We have $\mu(A)\leq\nu(A)$ for all $A\in R^{\ast}$.
  • (ii)
    For all $\varepsilon>0$, there exists a coupling $\pi_{\varepsilon}$ of $\mu$ and $\nu$ satisfying $\pi_{\varepsilon}(R)\geq 1-\varepsilon$.
However, the coupling counterpart statement (ii) cannot in general be improved.
The proof of Theorem 1.1, about which we now make a few remarks, is given in Section 2. First, we note that the negative part is shown, in Subsection 2.4, by directly constructing counterexamples in the setting of a simple $F_{\sigma}$ partial order on $[0,2]$ (Example 2.4) and the pre-order of eventual domination on $\mathbb{R}^{\mathbb{N}}$ (Example 2.5). Second, we note that the positive part involves a few steps, one of which is a general result in optimal transport that may be of independent interest. To state it, recall that the usual optimal transport theory (e.g., [34]) requires a cost function $c:S\times S\to\mathbb{R}$ which is lower semicontinuous. The following result shows that the Kantorovich dual problem is still well-behaved when $c$ lacks this regularity.
Theorem 1.2.
Suppose $S$ is a Polish space and that $c:S\times S\to[0,\infty)$ satisfies
  • (a)
    $c(x,x)=0$ for all $x\in S$,
  • (b)
    $c(x,z)\leq c(x,y)+c(y,z)$ for all $x,y,z\in S$, and
  • (c)
    there exists $\{c_{n}\}_{n\in\mathbb{N}}$ such that $c_{n}:S\times S\to[0,\infty)$ is lower semicontinuous for each $n\in\mathbb{N}$, and $c_{n}\downarrow c$ pointwise as $n\to\infty$.
Then, for all Borel probability measures $\mu,\nu$ on $S$ we have
(1.4) $\sup_{\begin{subarray}{c}\phi,\psi\in b\mathcal{B}(S)\\ \phi\oplus\psi\leq c\end{subarray}}\left(\int_{S}\phi\,\textnormal{d}\mu+\int_{S}\psi\,\textnormal{d}\nu\right)=\sup_{\begin{subarray}{c}\phi\in b\mathcal{B}(S)\\ \phi\ominus\phi\leq c\end{subarray}}\left(\int_{S}\phi\,\textnormal{d}\mu-\int_{S}\phi\,\textnormal{d}\nu\right)$
where $b\mathcal{B}(S)$ denotes the set of all bounded measurable functions from $S$ to $\mathbb{R}$.
Note, in particular, that this result applies to cost functions $c$ which are upper semicontinuous, since every upper semicontinuous function can be written as a non-increasing limit of continuous functions. However, we will apply the result to cost functions which are neither lower semicontinuous nor upper semicontinuous.
1.4. Future Work
First, there are questions arising from the connection we have identified between Thorisson’s working hypothesis and optimal transport for zero-one-valued cost functions. For instance, one may hope to understand cyclic monotonicity for such cost functions, hence [4, Theorem 1] establishing optimality conditions for the resulting optimal transport problem; this would imply some further structure of the various couplings statements appearing as statement (ii) above, as was previously done for equivalence relations in [17, Proposition 3.12]. For another example, one may hope to understand the existence of Monge solutions to the resulting optimal transport problem; this would imply that the couplings appearing in statement (ii) above can be achieved “without additional randomness,” as was previously done in [1] in the setting of measurable selections of random set-valued functions and in [16] in the setting of germ couplings of Brownian motions with drift.
Second, there are questions concerning applications of our general results to some problems of interest in probability theory. For instance, one may consider the problem of characterizing when two real-valued stochastic process $\{X_{t}\}_{t\geq 0}$ and $\{Y_{t}\}_{t\geq 0}$ can be coupled to satisfy eventual domination. We believe this problem is particularly interesting when $\{X_{t}\}_{t\geq 0}$ and $\{Y_{t}\}_{t\geq 0}$ are transient Markov processes, where a natural conjecture is that eventual domination should be closely related to stochastic domination of the induced laws on the common Martin boundary of the two processes. More concretely, one may aim to show that the existence of a coupling satisfying eventual domination for, say, two Lévy processes, is equivalent to some explicit property of their characteristic triplets or their Laplace exponents.
2. Proofs of Results
This section contains the proof of Theorem 1.1, as well as some intermediate results of independent interest. Throughout, we let $R\subseteq S\times S$ denote a reflexive, transitive, $F_{\sigma}$ relation, and we fix arbitrary Borel probability measures $\mu,\nu$ on $S$. The remaining subsections divide the proof into a few different steps: eliminating one of the two dual variables (Subsection 2.1), proving the existence of a dual maximizer (Subsection 2.2), converting this dual maximizer into a maximizing indicator function (Subsection 2.3), and constructing a counterexample to prove the possibility of non-existence of primal minimizers (Subsection 2.4).
2.1. Dual variable elimination
Note that $R$ being $F_{\sigma}$ implies that there exists a non-decreasing sequence $\{R_{n}\}_{n\in\mathbb{N}}$ of closed sets in $S\times S$ such that $R=\bigcup_{n\in\mathbb{N}}R_{n}$. Then, let us write $c=1-\mathbf{1}_{R}$ and $c_{n}=1-\mathbf{1}_{R_{n}}$ for each $n\in\mathbb{N}$. It follows that the hypotheses of Theorem 1.2 are satisfied, so, by combining with general results on Kantorovich duality (e.g., [25, Corollary 2.3.9]), we have
(2.1) $\sup_{\begin{subarray}{c}\phi\in b\mathcal{B}(S)\\ \phi\ominus\phi\leq c\end{subarray}}\left(\int_{S}\phi\,\textnormal{d}\mu-\int_{S}\phi\,\textnormal{d}\nu\right)=\inf_{\pi\in\Pi(\mu,\nu)}(1-\pi(R)).$
In the remainder of this subsection we prove Theorem 1.2, hence establishing (2.1). Note that this shows that the Kantorovich dual can in fact be taken to have one dual variable rather than two.
Proof of Theorem 1.2.
Let $\phi,\psi:S\to\mathbb{R}$ be arbitrary bounded, Borel measurable functions satisfying $\phi(x)+\psi(y)\leq c(x,y)$ for all $x,y\in S$, and write $\psi^{c}(x)\coloneqq\inf_{y\in S}(c(x,y)-\psi(y))$ for the $c$-transform of $\psi$. Now note that property (a) implies $\psi^{c}\leq-\psi$ pointwise, and that property (b) implies $\psi^{cc}=-\psi^{c}$. Moreover, by standard optimal transport considerations, we have
(2.2) $\int\phi\,d\mu+\int\psi\,d\nu\leq\int\psi^{c}\,d\mu+\int\psi\,d\nu\leq\int\psi^{c}\,d\mu-\int\psi^{c}\,d\nu.$
While it may appear that Theorem 1.2 follows immediately from (2.2), there is a problem in that $\psi^{c}$ need not be Borel measurable; indeed, if $\psi=\mathbf{\mathbf{1}}_{B}$ and $c=1-\mathbf{1}_{R}$, then $\psi^{c}=-\mathbf{\mathbf{1}}_{p((B\times S)\cap R)}$ where $p:S\times S\to S$ denotes the first-coordinate projection, in which case $p((B\times S)\cap R)$ analytic but possibly non-measurable.
We claim that $\psi^{c}$ is, in general, universally measurable, so that $\int\psi^{c}\,\textnormal{d}\mu,\int\psi^{c}\,\textnormal{d}\nu$ are well-defined via the $\mu$- and $\nu$-completions of $\mathcal{B}(S)$, respectively. To show this, note that the strict sub-level sets can be written as
$(\psi^{c})^{-1}((-\infty,\alpha))$ $=\left\{x\in S:\inf_{y\in S}(c(x,y)-\psi(y))<\alpha\right\}$
$=\{x\in S:\textnormal{there exists }y\in S\text{ such that }c(x,y)-\psi(y)<\alpha\}$
$=p(\{(x,y)\in S\times S:c(x,y)-\psi(y)<\alpha\}).$
As projections of Borel sets, these sub-level sets are universally measurable. (See [18, Theorem A1.2].)
Now let us show how to get around this measurability problem by directly constructing a Borel-measurable function $f$ with $f(x)-f(y)\leq c(x,y)$ pointwise, whose objective value on the right side of (1.4) is within $\varepsilon>0$ of the objective value of $(\phi,\psi)$ on the left side of (1.4), where $\varepsilon>0$ is fixed and arbitrary. First, without loss of generality, we may assume that $\psi\leq 0$, so that $\psi^{c}\geq 0$. Now let $\eta=\sum_{i=1}^{r}b_{i}\mathbf{1}_{E_{i}}$ be a (universally-measurable) simple function (with disjoint $E_{i}$) such that $\eta\leq\psi^{c}$ and $\int\eta\,\,\textnormal{d}\mu\geq\int\psi^{c}\,\,\textnormal{d}\mu-\varepsilon/2$. By the inner regularity of $\mu$, we may find (disjoint) compact sets $K_{i}\subseteq E_{i}$ such that $\mu(E_{i}\setminus K_{i})\leq\frac{\varepsilon}{2rb_{i}}$, so that, defining $g:=\sum_{i=1}^{r}b_{i}\mathbf{1}_{K_{i}},$
$\int g\,\,\textnormal{d}\mu=\sum_{i=1}^{r}b_{i}\mu(K_{i})\geq\sum_{i=1}^{r}b_{i}\left(\mu(E_{i})-\frac{\varepsilon}{2rb_{i}}\right)=\int\eta\,\,\textnormal{d}\mu-\frac{\varepsilon}{2}\geq\int\psi^{c}\,\,\textnormal{d}\mu-\varepsilon.$
If we can show that $g^{c}$ is Borel-measurable, we can finish by choosing $f=-g^{c}$ because $g^{c}(x)+g^{c}(y)=-g^{c}(x)+(-g^{c})^{c}(y)\leq c(x,y)$ and
$\int(-g^{c})\,\,\textnormal{d}\mu-\int(-g^{c})\,\,\textnormal{d}\nu$ $\geq\int g\,\,\textnormal{d}\mu-\int\psi^{c}\,\,\textnormal{d}\nu$
$\geq\int\psi^{c}\,\,\textnormal{d}\mu-\int\psi^{c}\,\,\textnormal{d}\nu-\varepsilon\geq\int\phi\,\,\textnormal{d}\mu-\int\psi\,\,\textnormal{d}\mu-\varepsilon$
where we have used $0\leq g\leq-g^{c}\leq-\psi^{cc}=\psi^{c}$.
To see that $g^{c}$ is Borel measurable, we again look at strict sub-level sets. And since $g\geq 0$, we know that $g^{c}\leq 0$, so we only need to look at $\alpha$ sub-level sets for $\alpha\leq 0$.
$(g^{c})^{-1}((-\infty,\alpha))$ $=\left\{x\in S:\inf_{y\in S}(c(x,y)-g(y))<\alpha\right\}$
$=\{x\in S:\textnormal{there exists }y\in S\text{ such that }c(x,y)-g(y)<\alpha\}$
$=p(\{(x,y)\in S\times S:c(x,y)-g(y)<\alpha\})$
$=p\left(\bigcup_{n=1}^{\infty}\{(x,y)\in S\times S:c_{n}(x,y)-g(y)<\alpha\}\right)$
For $y$ such that $g(y)=0$, i.e. $y\notin\bigcup_{i=1}^{r}K_{i}$, the condition becomes $c_{n}(x,y)<\alpha\leq 0$, which is impossible because $c_{n}\geq 0$. So we may assume that $y\in K_{1}\cup\cdots\cup K_{r}$. Then
$(g^{c})^{-1}((-\infty,\alpha))$ $=p\left(\bigcup_{n=1}^{\infty}\bigcup_{i=1}^{r}\{(x,y)\in S\times K_{i}:c_{n}(x,y)-b_{i}<\alpha\}\right)$
$=\bigcup_{n=1}^{\infty}\bigcup_{i=1}^{r}p(\{(x,y)\in S\times K_{i}:c_{n}(x,y)<\alpha+b_{i}\}).$
Finally, we claim that each of these projections is closed and hence Borel-measurable. Suppose $x_{m}\in p(\{(x,y)\in S\times K_{i}:c_{n}(x,y)<\alpha+b_{i}\})$ with $x_{m}\to x\in S$; we want to show that $x$ also lies in the projection. There exist $y_{m}\in K_{i}$ such that $c_{n}(x_{m},y_{m})<\alpha+b_{i}$, and by the compactness of $K_{i}$, by passing to a subsequence, we may assume that $y_{m}\to y\in K_{i}$. Then $y$ witnesses the membership of $x$ in the projection, as
$c_{n}(x,y)\leq\liminf_{m\to\infty}c_{n}(x_{m},y_{m})<\alpha+b_{i},$
where we have used the lower semicontinuity of $c_{n}$. So these projections are closed, and we conclude that $g^{c}$ is Borel-measurable. ∎
Remark 2.1.
Discussions about the Borel measurability of $c$-transformed functions appear already in most standard introductions to optimal transport (e.g., [34, Remark 5.5 and p. 69]). Our result required slightly more work, since our cost function $c$ need not even be lower semi-continuous.
2.2. Existence of dual maximizer
Our next step is to show that the Kanotorovich dual problem, i.e., the left side of (2.1), admits a maximizer for $c=1-\mathbf{1}_{R}$.
Remark 2.2.
The analysis in this subsection does not require any hypothesis on the function $c=1-\mathbf{1}_{R}$ beyond being uniformly bounded. So, the same argument applies to the two-variable appearing on the left side of (1.4).
The idea is to take a weak limit of approximate maximizers to get a maximizer. Consider the Hilbert space $\mathcal{H}:=L^{2}(\rho)\times L^{2}(\rho)$, where $\rho:=(\mu+\nu)/2$; the choice of $L^{2}(\rho)\times L^{2}(\rho)$ rather than $L^{2}(\mu)\times L^{2}(\nu)$ will be useful later. Let $\phi_{n}\in b\mathcal{B}(S)$ (with $\phi_{n}\ominus\phi_{n}\leq c$) be a sequence of $\nicefrac{{1}}{{n}}$-approximate maximizers for the Kantorovich dual problem. The sequence $\{(\phi_{n},-\phi_{n})\}_{n\in\mathbb{N}}$ is norm-bounded in $\mathcal{H}$ since it is pointwise uniformly bounded, so the Banach-Alaoglu theorem guarantees that there exist some $\{n_{k}\}_{k\in\mathbb{N}}$ and $(\phi_{\infty},-\phi_{\infty})\in\mathcal{H}$ such that we have $(\phi_{n_{k}},-\phi_{n_{k}})\to(\phi_{\infty},-\phi_{\infty})$ weakly.
However, notice that the functions $\phi_{\infty},-\phi_{\infty}$ are only defined up to null sets, so we cannot hope to show $\phi_{\infty}\ominus\phi_{\infty}\leq c$, which is a pointwise constraint. To obtain a pointwise cost inequality, we need to show that $\phi_{\infty}$ admits a modification $\phi^{\prime}_{\infty}$ so that $(\phi^{\prime}_{\infty},-\phi^{\prime}_{\infty})$ together satisfy the pointwise constraint. The key to this is the following intermediate result, which is essentially equivalent to main result of [14] which answers a measure-theoretic question initially posed by Arveson [2]; we believe it may be of interest in similar works on optimal transport in the future.
Lemma 2.3.
For $f,g:S\to\mathbb{R}$ and $\ell:S\times S\to[0,\infty)$ any Borel-measurable functions, the following are equivalent:
  1. (a)
    For all $\pi\in\Pi(\mu,\nu)$ we have $f(x)+g(y)\leq\ell(x,y)$ holding for $\pi$-almost all $(x,y)\in S\times S$.
  2. (b)
    There exist Borel measurable functions $f^{\prime},g^{\prime}:S\to\mathbb{R}$ such that $f^{\prime}(x)=f(x)$ holds for $\mu$-almost all $x\in S$, $g^{\prime}(y)=g(y)$ holds $\nu$-almost all $y\in S$, and $f^{\prime}(x)+g^{\prime}(y)\leq\ell(x,y)$ holds for all $x,y\in S$.
Proof.
The claim (b) implies (a) is immediate, so we only need to prove that (a) implies (b). Let $E:=\{(x,y)\in S\times S:f(x)+g(y)>\ell(x,y)\}$, and note by assumption that we have $\pi(E)=0$ for all $\pi\in\Pi(\mu,\nu)$. By [14, Corollary, p. 500] this implies that there exist Borel sets $N_{f},N_{g}\subseteq S$ with $\mu(N_{f})=\nu(N_{g})=0$ such that $E\subseteq(N_{f}\times S)\cup(S\times N_{g})$. Now we set
[…]
and similar for $g^{\prime}$. Then $f^{\prime}$ and $g^{\prime}$ are as desired. ∎
To apply Lemma 2.3, we need to check that $(\phi_{\infty},-\phi_{\infty})$ satisfies the cost inequality $\pi$-a.s. for any $\pi\in\Pi(\rho,\rho)$. Define
$h(x,y)$ $:=1-\mathbf{1}_{R}(x,y)-\phi_{\infty}(x)+\phi_{\infty}(y),\qquad\textnormal{ and}$
$h_{k}(x,y)$ $:=1-\mathbf{1}_{R}(x,y)-\phi_{n_{k}}(x)+\phi_{n_{k}}(y)$
for all $k\in\mathbb{N}$, and observe that $\phi_{n_{k}}\ominus\phi_{n_{k}}\leq c$ implies $h_{n_{k}}\geq 0$. Now take an arbitrary $\pi\in\Pi(\rho,\rho)$, and notice that that $h_{n_{k}}\to h$ weakly in $L^{2}(\pi)$. Thus, by the definition of weak convergence, we have:
$0\leq\lim_{k\to\infty}\int_{S\times S}h_{k}\mathbf{1}_{h\leq 0}\,\textnormal{d}\pi=\int_{S\times S}h\mathbf{1}_{h\leq 0}\,\textnormal{d}\pi\leq 0,$
so that $\pi(\{h<0\})=0$.
Lemma 2.3 now shows that we can adjust $\phi_{\infty}$$\phi_{\infty}$ by setting them to be 0 on $\rho$-null sets $N_{\phi},N_{-\phi}$ to obtain Borel-measurable versions $\phi_{\infty}^{\prime},(-\phi_{\infty})^{\prime}$ which satisfy the cost inequality pointwise; by making the modification using $N_{\phi}\cup N_{-\phi}$ instead for both functions, we can guarantee that $(-\phi_{\infty})^{\prime}=-\phi_{\infty}^{\prime}$ pointwise. And since the set $N_{\phi}\cup N_{-\phi}$ is $\rho$-null and hence both $\mu$-null and $\nu$-null (this is where we use the choice of $L^{2}(\rho)\times L^{2}(\rho)$ rather than $L^{2}(\mu)\times L^{2}(\nu)$), we can use weak convergence to get
$\int\phi_{\infty}^{\prime}\,\,\textnormal{d}\mu-\int\phi_{\infty}^{\prime}\,\,\textnormal{d}\nu$ $=\int\phi_{\infty}\,\,\textnormal{d}\mu-\int\phi_{\infty}\,\,\textnormal{d}\nu$
$=\int\phi_{\infty}\left(\frac{\,\textnormal{d}\mu}{\,\textnormal{d}\rho}-\frac{\,\textnormal{d}\nu}{\,\textnormal{d}\rho}\right)\,\,\textnormal{d}\rho$
$=\lim_{k\to\infty}\int\phi_{n_{k}}\left(\frac{\,\textnormal{d}\mu}{\,\textnormal{d}\rho}-\frac{\,\textnormal{d}\nu}{\,\textnormal{d}\rho}\right)\,\,\textnormal{d}\rho$
where we used $0\leq\frac{\,\textnormal{d}\mu}{\,\textnormal{d}\rho},\frac{\,\textnormal{d}\nu}{\,\textnormal{d}\rho}\leq 2$ hence $\frac{\,\textnormal{d}\mu}{\,\textnormal{d}\rho}-\frac{\,\textnormal{d}\nu}{\,\textnormal{d}\rho}\in L^{2}(\rho)$. Then
$\int\phi_{\infty}^{\prime}\,\,\textnormal{d}\mu-\int\phi_{\infty}^{\prime}\,\,\textnormal{d}\nu$ $=\lim_{k\to\infty}\int\phi_{n_{k}}\,\,\textnormal{d}\mu-\int\phi_{n_{k}}\,\,\textnormal{d}\nu$
$\geq\lim_{k\to\infty}\sup_{\begin{subarray}{c}\phi\in b\mathcal{B}(S)\\ \phi\ominus\phi\leq c\end{subarray}}\int\phi\,\,\textnormal{d}\mu-\int\phi\,\,\textnormal{d}\nu-\frac{1}{n_{k}}$
$=\sup_{\begin{subarray}{c}\phi\in b\mathcal{B}(S)\\ \phi\ominus\phi\leq c\end{subarray}}\int\phi\,\,\textnormal{d}\mu-\int\phi\,\,\textnormal{d}\nu.$
2.3. Conversion to indicator maximizer
As the last step of the affirmative direction, we show that the Kantorovich dual problem, i.e., the left side of (2.1), can be taken over pairs of indicator functions rather than over pairs of bounded measurable functions.
To do this, let $\phi_{\infty}^{\prime}:S\to\mathbb{R}$ denote the dual maximizer from the previous subsection. Recall from the usual “layer cake” decomposition that we can write
$(\phi_{\infty}^{\prime}(x),-\phi_{\infty}^{\prime}(y))=\int_{0}^{1}(\mathbf{1}_{A_{t}}(x),-\mathbf{1}_{A_{t}}(y))\,\,\textnormal{d}t,$
where we have defined $\mathbf{1}_{A_{t}}:=\{x\in S:t\leq\phi_{\infty}^{\prime}(x)\leq 1\}$ for all $0\leq t\leq 1$.
To conclude, we claim that $A_{t}\in R^{\ast}$ holds for all $0\leq t\leq 1$. Indeed, if $(x,y)\in S\times S$ satisfies $(x,y)\notin R$ or $\phi_{\infty}^{\prime}(x)<t$, then we of course have $\mathbf{1}_{A_{t}}(x)-\mathbf{1}_{A_{t}}(y)\leq 1-\mathbf{1}_{R}(x,y)$; otherwise, we have $(x,y)\in R$ and $t\leq\phi_{\infty}^{\prime}(x)$, so we conclude $\phi_{\infty}^{\prime}(y)\leq-\phi_{\infty}^{\prime}(x)\leq-t$ hence $\mathbf{1}_{A_{t}}(x)-\mathbf{1}_{A_{t}}(y)=1-1=0=1-\mathbf{1}_{R}(x,y)$.
Now note that by Fubini-Tonelli, we have:
$\int_{S}\phi_{\infty}^{\prime}\,\textnormal{d}\mu-\int_{S}\phi_{\infty}^{\prime}\,\textnormal{d}\nu$ $=\int_{S}\int_{0}^{1}\mathbf{1}_{A_{t}}(x)\,\textnormal{d}t\,\textnormal{d}\mu(x)-\int_{S}\int_{0}^{1}\mathbf{1}_{A_{t}}(y)\,\textnormal{d}t\,\textnormal{d}\nu(y)$
$=\int_{0}^{1}\left(\int_{S}\mathbf{1}_{A_{t}}(x)\,\textnormal{d}\mu(x)-\int_{S}\mathbf{1}_{A_{t}}(y)\,\textnormal{d}\nu(y)\right)\,\textnormal{d}t$
$=\int_{0}^{1}(\mu(A_{t})-\nu(A_{t}))\,\textnormal{d}t$
$\leq\sup_{A\in R^{*}}(\mu(A)-\nu(A)),$
establishing the affirmative part of Theorem 1.1.
2.4. Non-existence of primal minimizers
The only remaining step of the proof is to show that the primal problem, i.e. the right side of (2.1), can fail to admit minimizers. Towards this end, we construct two counterexamples; Example 2.4 is elementary, and Example 2.5 is an embedding of Example 2.4 into a more complicated setting of interest.
11221122xxyy
Figure 2.1. Counterexample to Strassen’s theorem for $F_{\sigma}$ partial orders.
Example 2.4.
Let $S=[0,2]$, and let $\mu$ and $\nu$ denote the Lebesgue measure on the subintervals $[0,1]$ and $[1,2]$, respectively. Then consider the following relation, which can also be visualized with the help of Figure 2.1:
$R:=\{(x,y)\in S\times S:x=y\textnormal{ or }y>x+1\}.$
Observe that $R$ is an $F_{\sigma}$ set, since we can write $R=\bigcup_{n\in\mathbb{N}}R_{n}$, where
(2.3) $R_{n}:=\left\{(x,y)\in S\times S:x=y\textnormal{ or }y\geq x+\frac{1}{n}\right\}.$
Also, $R$ is a partial order, and in particular it is reflexive and transitive.
First, we claim that the primal value is zero. To see this, fix $0<\varepsilon<1$ and define the map $T_{\varepsilon}(x):[0,1]\to[1,2]$ via $T(x):=1+[x-\varepsilon]$, where $\alpha]$ denotes the fractional part of a real number $\alpha\in\mathbb{R}$; then define $\pi_{\varepsilon}\in\Pi(\mu,\nu)$ as the pushforward $\pi_{\varepsilon}:=(\textnormal{id},T_{\varepsilon})_{\#}\mu$. Note $\pi_{\varepsilon}(R)=1-\varepsilon$ for all $0<\varepsilon<1$, so taking $\varepsilon\to 0$ shows that the primal value is zero.
Second, we claim that there is no coupling achieving a primal value of zero. To see this, assume towards a contradiction that $\pi$ is a coupling of $\mu,\nu$ with $\pi(R)=1$. Then let $X$ and $Y$ be random variables defined on a common probability space, with probability $\mathbb{P}$ and expectation $\mathbb{E}$, such that the joint distribution of $(X,Y)$ is $\pi$. Since the distributions $\mu,\nu$ of $X,Y$ are mutually singular, we have $\mathbb{P}(X=Y)=0$; By combining this with $\mathbb{P}((X,Y)\in R)=1$, we conclude $\mathbb{P}(Y>X+1)=1$. Taking expectations yields $\mathbb{E}[Y]>\mathbb{E}[X]+1$, but this contradicts the marginal constraints which imply $\mathbb{E}[X]=\int_{0}^{1}x\,\textnormal{d}x=\nicefrac{{1}}{{2}}$ and $\mathbb{E}[Y]=\int_{1}^{2}x\,\textnormal{d}x=\nicefrac{{3}}{{2}}$. This shows that there is no coupling achieving the primal value.
Example 2.4 is counterexample to the claim “Thorisson’s working hypothesis holds for $F_{\sigma}$ partial orders,” although only insofar as the primal problem may fail to admit minimizers. In light of analogous work on equivalence relations [17, Proposition 3.18], it is natural to wonder whether the weaker claim “Thorisson’s working hypothesis holds for partial orders that can be written as a countable increasing union of closed partial orders” is true; note, however, that Example 2.4 is a counterexample to this statement as well, since the closed sets in (2.3) are indeed partial orders. This discrepancy highlights that there is further structure in equivalence relations, compared to in partial orders, which allows for a more complete duality theory.
Although the partial order $([0,2],R)$ from Example 2.4 may seem artificial, we note that it can be embedded into many other orders of interest, hence establishing counterexamples in other settings. Of particular interest is the pre-order of eventual domination from Subsection 1.4, which we denote $(\mathbb{R}^{\mathbb{N}},\leq^{\ast})$.
Example 2.5.
Let $S=\mathbb{R}^{\mathbb{N}}$, let $\leq^{\ast}$ denote the pre-order of eventual domination, and let $([0,2],R)$ be the partially ordered set defined in Example 2.4. Our goal is to construct a function $\phi$ such that $\phi:([0,2],R)\to(\mathbb{R}^{\mathbb{N}},\leq^{\ast})$ is an order embedding, meaning we have $(x,y)\in R$ if and only if $\phi(x)\leq^{\ast}\phi(y)$. Indeed, we may set
$\phi(z):=\begin{cases}a(z)&\textnormal{ if }0\leq z\leq 1\\ b(z-1)&\textnormal{ if }1<z\leq 2\\ \end{cases},\qquad\textnormal{where}$
[…]
It can be checked by casework that $\phi$ is an order embedding. In order to transfer Example 2.4 to the present setting, we let $\mu$ and $\nu$ be the Lebesgue measures on the intervals $[0,1]$ and $[1,2]$, respectively, and then push these forward to $\tilde{\mu}:=\phi_{\#}\mu$ and $\tilde{\nu}:=\phi_{\#}\nu$ on $\mathbb{R}^{\mathbb{N}}$.
First, we claim that the primal problem has value zero. To see this, note that for any $\pi\in\Pi(\mu,\nu)$ we have
$\int_{S\times S}(1-\mathbf{1}_{R_{\leq^{\ast}}})\,\textnormal{d}\left((\phi,\phi)_{\#}\pi\right)=\int_{[0,2]\times[0,2]}(1-\mathbf{1}_{R})\,\textnormal{d}\pi.$
Hence, an $\varepsilon$-optimal coupling for the primal problem in $([0,2],R)$ leads to an $\varepsilon$-optimal coupling for the primal problem in $(S,\leq^{\ast})$, and we conclude that the optimal value is zero.
[…]
$\int_{[0,2]\times[0,2]}(1-\mathbf{1}_{R})\,\textnormal{d}\pi=\int_{S\times S}(1-\mathbf{1}_{R_{\leq^{\ast}}})\,\textnormal{d}\tilde{\pi}=0.$
This contradicts Example 2.4 since no such $\pi$ can exist.